• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Puckett

    Log entry time 08:03:31 on December20,2008

    Entry number 254129

    keyword=Shift summary

    Owl shift 12/20/2008:

    Took over from swing shift on the tail end of a ~30 minute beam down. We prepared to move the target to BeO for the standard checkout procedure following 1+ hr downtime, and then MCC called to inform us they were ready to restore beam. So we did not move the target and instead resumed production data taking on 3He. We started with 2 uA while MCC steered the beam to the correct BPMA and BPMB values, and then ramped up to 12 uA without interrupting the CODA run. When the beam came back, several BigBite HV channels tripped, these were restored without difficulty.

    Run Summary:

    Run #, start, stop, charge
    4618, 00:18, 01:14, .0354 (u3)
    4619, 01:16, 02:09, .03573 (u3)
    4620, 02:10, 03:06, .03542 (u3) (started new HAPPEX run 30600)
    4621, 03:08, 04:03, .03528 (u3)
    4622, 04:04, 05:08, .03505 (u3)
    4623, 05:08, 06:06, .03519 (u3)
    4624, 06:07, 07:03, .03545 (u3)
    4625, 07:04, 08:00, .03545 (u3)